OpenStudy (nuccioreggie):

What is the value of z in the equation 2(4z − 9 − 7) = 166 − 46? 19 21 26 34

OpenStudy (whovianchick):

2 (4z-9-7) = 166-46 2(4z-16) = 120 (combine like terms) 8z-32 = 120 (distribute) 8z-32=120 (combine like terms) 8z=152 (add 32 to both sides) z=19 (divide both sides by 8) A: 19

OpenStudy (whovianchick):

am I making sense on all of these?

OpenStudy (nuccioreggie):

yes you are

OpenStudy (whovianchick):

good just making sure since we're moving fast :)